OpenStudy (anonymous):

A dragster accelerates at 9m/s^2 from a stationary start for 7 seconds. How far does it travel in 7 seconds?

OpenStudy (anonymous):

dx= (initial velocity) (time) + 1/2 ( acceleration) (time^2) initial velocity= 0 time= 7 secs acceleration= 9m/s^2 dx= (0) ( 7 secs) + 1/2 (9 m/s^2) (7^2) = 0+ 1/2 (9 m/s^2) (49) = 1/2 (441) = 220.5 m Lol. Not sure.

OpenStudy (anonymous):

That's it. Thanks man.